### v1

# Wichitrnithed, August 28, 2025 GPU-acceleration of the Discontinuous Galerkin Shallow Water Equations Solver (DG-SWEM) using CUDA and OpenACC

Abstract
This paper presents a porting of DG-SWEM, a discontinuous Galerkin solver for coastal ocean circulation, and in particular storm surge, to GPU using two separate approaches: CUDA Fortran and OpenACC. Time-explicit discontinuous Galerkin methods have been shown to exhibit a large amount of data parallelism due to the loose coupling between elements, and thus are naturally mapped to the GPU architecture. For each porting approach, we discuss the code design, ease of programming, and performance when running on realistic use cases. Specifically for the OpenACC version, we also aim to preserve maintainability within the same codebase through using Unified Memory. We test the codes on NVIDIA’s Grace Hopper chip and compare the GPU performance on each node to the MPI version on a single CPU node (144 cores).
1 Introduction
One of the primary causes of flooding on the Texas and US Gulf and Atlantic coasts is storm surge: the inundation of coastal regions by water in the ocean that is driven inland by hurricanes or tropical storms. This phenomenon is extremely devastating to life and property of those nearshore as well as coastal ecosystems. In mitigating damage and guiding relief and emergency efforts, computer simulation of storm surge is critical in both operational forecasting as well as in hindcasting past storms for validation, engineering planning, and design studies. Some well-established and commonly used models include, for storm surge, the ADvanced CIRCulation (ADCIRC) model [Pringle et al. (2020), Luettich et al. (1992)], the Sea, Lake, and Overland Surges from Hurricanes (SLOSH) [Jelesnianski (1992)], and Delft3D [Veeramony et al. (2017)]; they have all been extensively validated for past hurricane events.
In this paper, the focus is on a model closely related to ADCIRC, the Discontinuous Galerkin Shallow Water Equation Model (DG-SWEM) [Kubatko et al. (2006), Dawson et al. (2011), Wichitrnithed et al. (2024)] and aims to improve the main criticism of this model, i.e., the computational cost. The formulation of ADCIRC is based on the two-dimensional shallow water equations (SWE) which are obtained by depth-averaging the incompressible Navier-Stokes equations. The advantage of using finite elements is the ability to model complex coastal domains through unstructured meshes. At the same time, a straightforward usage of the continuous Bubnov-Galerkin (CG) method results in spurious oscillations for advection-dominated flows. ADCIRC handles this by using a reformulation of the continuity equation in the SWE called the Generalized Wave Continuity Equation (GWCE), but as a consequence of this and the CG methodology, local mass conservation is no longer guaranteed. Furthermore, in practice, it still necessary to turn the advective terms in the GWCE off for some cases to achieve numerical stability.
One critical type of flooding that presents difficulties for ADCIRC is compound flooding which arises when multiple water sources such as storm surge and rainfall with resulting river discharge occur simultaneously. In practice, ADCIRC has encountered instability when rainfall is added to the source term of the GWCE and its subsequent interaction with wetting and drying scheme. DG-SWEM has recently [Wichitrnithed et al. (2024)] been shown to be able to address many of these problems by using the discontinuous Galerkin (DG) finite element formulation on the primitive SWE [Dawson et al. (2011), Dawson and Proft (2002)]. The DG method can be viewed as a synthesis of the Galerkin method and the finite volume method, whereby the solution is approximated using polynomial basis functions local to each finite element (cell) which guarantees local conservation. Solutions need not have a single value at each node (vertex) and coupling between adjacent elements is done through a numerical flux. It has been widely applied to hyperbolic conservation laws in various domains such as electromagnetics, acoustics, plasma physics, and shallow water equations. We refer to [Hesthaven and Warburton (2008), Cockburn (1998)] for the history and details of the DG method.
A primary disadvantage of the DG method which has confined DG-SWEM to mainly research and not operational forecasting like ADCIRC is the increased computational cost associated with the computations required from the higher number of global degrees of freedom compared to the continuous formulation. At the same time, the localized nature of computing the solution for each element independently generates a high degree of data parallelism, and several GPU implementations of the DG method have been using different approaches. Early implementations using NVIDIA’s CUDA include [Klöckner et al. (2009), Siebenborn et al. (2012), Fuhry (2013), Brodtkorb et al. (2012)]. More recent implementations using different application programming interfaces (API) can be seen in [Abdi et al. (2019), Kirby and Mavriplis (2020)] which use OCCA, and [Xia et al. (2014), Dai et al. (2022)] which use OpenACC. This paper presents and compares the porting of DG-SWEM to GPU using two separate approaches: CUDA and OpenACC. The aim of the latter approach is to explore a directive-based programming model which requires less integration and maintenance efforts, as well as to take advantage of the new NVIDIA Grace-Hopper architecture [Singer (2024)] which provides a more efficient CPU-GPU data transfer. Lastly, we mention the recent open-source FEniCSx SWE solverby Dawson et al. in [Dawson et al. (2024)] and GPU accelerated by Pachev in [Pachev (2024)], where implicit schemes are considered.
In the following, we introduce, discuss, and validate the proposed GPU acceleration. In section 2, we introduce the governing model, i.e., the SWE and our DG discretization. Next, in Sections 3 and 4, we discuss the acceleration approach and the hardware setup we utilize. In 6, we present numerical studies of our methodology including benchmark tests and large scale hurricanes that include compound flood effects. Finally, in Sections 7 and 8, we discuss our results, draw conclusions, and discuss potential future research directions.
2 Model
2.1 Shallow water equations
The governing model in DG-SWEM is the two-dimensional shallow water equations which consists of mass and momentum conservation in each direction. The problem can be visualized as in Figure 1 and formulated as follows [Tan (1992)]:
where:
is the free surface elevation (positive upwards from geoid),
the bathymetry (positive downwards from the geoid),
the total water column,
the depth-averaged velocity field,
the bottom friction factor,
the mass source term such as rain,
the relevant sources which induce flow, including, e.g., Coriolis force, tidal potential forces, wind stresses, and wave radiation stresses, and
the computational domain, e.g., the coastal ocean, and its boundary is identified by three distinctive sections .
2.2 The Discontinuous Galerkin Formulation
The particular DG formulation and solution method used in this work is based on previous works [Kubatko et al. (2006), Dawson et al. (2011)]. Hence, we do not include a comprehensive introduction to the details of the formulation but rather a brief overview to make this presentation sufficiently self-contained. This description is an updated version from our recent work [Wichitrnithed et al. (2024)].
First, we define the vector of state variables to be solved, , the source vector and the flux matrix as:
we can write these three equations a conservative form
As is standard in the discretization of (2) with a finite element formulation, we partition the domain into (triangular) non-overlapping elements,
and the corresponding weak formulation is obtained by multiplying each equation by a test function and integrating over each element and subsequently performing an integrating by parts to the term with the spatial derivative:
where we use standard inner product notation for integration over elements, i.e., and element edges . represents the numerical flux, e.g., Local Lax-Friedrichs or Roe flux. The role of the numerical flux is to couple adjacent, discontinuous elements and correctly capture the flow characteristics across the edges; this is necessary for the solution to be stable [Hesthaven and Warburton (2008), Chapter 2]. Thus, aside from calculating (which depends on two neighboring elements), the computation can proceed on each element independently.
In accordance with the discontinuous Galerkin method, we next approximate the solution in space as using a set of basis functions. We opt to use the orthogonal Dubiner basis [Dubiner (1991)] as:
where are the modal degrees of freedom. The indices indicate the order of the polynomial: is of order . This basis is chosen as it produces a diagonal mass matrix, and higher-order approximations can be obtained by simply adding more terms. Note that while DG-SWEM supports higher-order polynomial orders, this work implements only the linear case, i.e. the first three basis functions are used. In the context of simulating storm surges and similar to ADCIRC, we found that using linear (or even constant) approximation is sufficient. Substituting (4) into the weak formulation (3) above, and requiring that it holds for all basis functions reduces the problem to a system of ODEs in the form:
where the terms from (3) have been collected into the RHS . Numerical integration of each term is performed using Gauss-Lobatto quadrature.
To integrate in time, we use an optimized version of the explicit Strong Stability Preserving Runge-Kutta scheme (SSPRK) based on [Shu and Osher (1988)] to discretize (5) in time and compute the difference approximation of , at the next timestep. This scheme has been specifically optimized with respect to DG spatial discretizations to enhance the CFL stability regions and therefore increase the possible time step size for a given order and stage [Kubatko et al. (2014)]. This is also combined with an optional slope limiter to suppress oscillations. Finally, to output the state variables, we convert the modal representation to a nodal representation by averaging the neighboring elements of each node.
In terms of execution, DG-SWEM supports parallel execution on CPUs through domain decomposition with METIS [Karypis and Kumar (1997)] using the Message-Passing Interface (MPI). As the scheme is explicit in both space and time, most computations are local to each element, except the calculation of the numerical flux that is shared between two neighboring elements. We use the standard ghost-cell technique to transfer state variables of elements at the boundary of each domain. An evaluation of the parallel efficiency of the code is reported in [Kubatko et al. (2009)].
3 GPU Programming approach
As an alternative to CPU parallelism, graphics processing units (GPUs) can be viewed as a multithreaded vector processor [Hennessy and Patterson (2018)], where each unit executes the same instructions on multiple data (SIMD), but individual lanes can take different paths. To take advantage of a large number of such processors, the algorithm must exhibit a high degree of data parallelism, where multiple elements can be processed independently at the same time. This is particularly applicable to explicit DG methods due to the inherently local computation of the elements.
In the past, as GPUs have originally been designed for graphics programming, the available application programming interface (API) such as DirectX and OpenGL imposed limitations of using GPUs for general-purpose computing (GPGPU) [Kirk and Hwu (2010)]. In 2007, NVIDIA developed CUDA [NVIDIA (2024)] which provides a more generic API to GPU instructions. This significantly extended the capabilities of NVIDIA GPUs to scientific and engineering codes: the programmer gained control of individual GPU threads, synchronization, and shared memory. Using the NVIDIA HPC compiler stack (a successor of PGI compilers), CUDA is supported in C, C++, and Fortran [NVIDIA (2020)]; CUDA code can be written in the same source files and compiled with appropriate flags. While the low-level nature of CUDA allows more opportunities for control and optimization by the programmer, it typically results in having to maintain a separate code base from the CPU code.
Another approach to GPU programming is to use OpenACC [OpenACC-Standard.org (2021)] and OpenMP [Openmp.org (2021)]
to write directives instead of manually managing individual GPU threads. This approach allows the programmer to insert high-level statements that generate GPU instructions and appear as comments in the source code (similar to OpenMP loop constructs, for instance). This can sometimes result in a single code base — only the compilation flags such as -acc need to be turned on to enable the GPU code generation. Of course, the programmer is still required to sufficiently expose the parallelism in the code, which often results in modifications of the CPU code as well.
In this work, we port DG-SWEM using each approach and compare their implementations and performance. For the directive-based approach, we choose to use OpenACC as it currently has better support on the NVHPC compilers.
4 Experimental setup
The primary hardware used in this work is the recent Vista Supercomputer [Singer (2024)] at the Texas Advanced Computing Center (TACC). Vista uses the ARM-based NVIDIA GH200 Grace Hopper architecture. Each GPU node consists of an NVIDIA H100 GPU connected to an NVIDIA Grace CPU directly through high-bandwidth NVLink-C2C interconnect. The Grace CPU contains 72 Arm Neoverse V2 cores and 120 GiB of LPDDR5X memory, while the Hopper GPU contains 96 GB of HBM3 memory. Each CPU node in the CPU-only subsystem contains 144 Grace cores on two sockets with a total memory of 237 GB LPDDR5. Different nodes are connected in a fat-tree topology, with GG nodes through an InfiniBand 200 GB/s and GH nodes an InfiniBand 400 GB/s. Specifications for these nodes are listed in Table 1 and 2.
The NVLink interconnect provides a faster data transfer between CPU and GPU. Previous generations of GPUs connect to CPUs through PCIe and trigger data migration on page faults. GH200 uses a hardware-based approach to achieve the same functionality by sharing the same page table and virtual address space which results in less page faults. In other words, CPU and GPU memories are now treated as NUMA nodes [Fusco et al. (2024)].
| CPU | NVIDIA Grace, 72 cores @ 3.1 GHz |
| CPU memory | 120 GB LPDDR5X |
| GPU | NVIDIA H100 |
| No. of SMs | 132 |
| FP64 | 34 TFLOPs/s |
| FP32 | 67 TFLOPs/s |
| GPU memory | 96 GB HBM3 |
| GPU memory bandwidth | 4 TB/s |
| NVLink-C2C CPU-GPU bandwith | 900 GB/s |
| CPU | NVIDIA Grace, 144 cores @ 3.4 GHz |
| FP64 | 7 TFLOPs/s |
| CPU memory | 237 GB LPDDR5 |
| CPU memory bandwidth | 850 GiB/s |
The original CPU code was compiled using the Intel oneAPI compiler stack. This work uses the NVIDIA SDK compiler stack (nvfortran, nvc) for both CPU and GPU versions, with the latter also using CUDA Fortran. Compilation flags for the CPU and GPU codes are listed in Table 3 . On Vista, the NVHPC version is 24.7 while on the local setup it is 23.11.
| Version | Flags |
| CUDA | nvfortran -r4 -O3 -cuda -Mextend -Mpreprocess - |
| OpenACC | nvfortran -r4 -O3 -cuda -acc -Mextend -Mpreprocess -gpu=unified |
| CPU | mpif90 -r4 -O3 -Mextend -Mpreprocess |
5 Implementation
This section details our implementation for both CUDA Fortran and OpenACC for different aspects of the code. In the Kernels section, a general parallelization is outlined, then details specific to either approaches are explained.
5.1 Data management
A significant part of porting code to GPUs is managing data movement. As a CPU (host) and a GPU (device) have separate physical memories, data must be transferred back and forth to perform the respective calculations. Even if some arrays are only used in device calculations, they still need to be initialized or read from files through the host, and we would have to declare device copies of those arrays. For example, an existing array in the code
real, allocatable :: A(:)
will need an additional copy on the device
real, allocatable, device :: A_d(:)
and copied to and from the device as necessary:
A_d = A
! Use A_d on the GPU
A = A_d
To avoid having to do this for each array in the code, we use CUDA’s unified and constant memory types, explained below.
5.1.1 Unified memory
Unified Virtual Addressing (UVA) was introduced in CUDA 6.0. This allows GPU kernels to access data on the host without needing to explicitly copy the data to device. Data migration is automatically performed once a page fault is triggered on the device. One way to achieve this is through the managed attribute. This way, we can reuse the existing arrays in the code and only modify their declarations, e.g.
real, allocatable, managed :: A(:,:)
! Use A on the device
istat = cudaDeviceSynchronize()
! Use A on the host
For newer systems that support Heterogenous Memory Management (HMM), the declaration can be completely eliminated using the compilation flag -gpu=unified. This significantly simplifies programming efforts and enhances maintainability. In both approaches, a synchronization is needed for correctness before the CPU can access data, since CUDA subroutines immediately return control to the host.
In this work, the CUDA Fortran approach uses the managed declarations to make explicit to the programmer which arrays will be used on the GPU, while the OpenACC code uses -gpu=unified to avoid any data declarations.
5.1.2 Constant memory
The standard way of passing scalar variables to CUDA kernels is by value, i.e. through the Fortran value attribute in the dummy argument declaration. When there are a lot of variables, such as constants, this increases the programmer’s efforts to pass them correctly. To avoid this, we use CUDA’s constant attribute for constant variables. They are declared and initialized in the host global module, and reside as read-only variables on the GPU constant memory partition which a size limit of 64 KB. Additionally, the constant memory has faster access time than global memory when all threads access the same element — a common pattern for many fixed-size arrays like quadrature weights. In Table 4, we summarize the important arrays whose attributes are modified, some of which are referred to in the kernel outline in the next section.
| Array | Meaning | Attribute |
| ZE[k,e,irk] | mode of elevation solution at element on RK stage irk | managed |
| QX[k,e,irk] | mode of x-velocity solution at element on RK stage irk | managed |
| QY[k,e,irk] | mode of y-velocity solution at element on RK stage irk | managed |
| RHS_ZE[k,e,irk] | Contribution to the mode of at element on RK stage irk | managed |
| RHS_QX[k,e,irk] | Contribution to the mode of at element on RK stage irk | managed |
| RHS_QY[k,e,irk] | Contribution to the mode of at element on RK stage irk | managed |
| PHI_AREA[k,i] | basis evaluated at area quadrature point i | constant |
| PHI_EDGE[k,i] | basis evaluated at edge quadrature point i | constant |
| XFAC[k,i,e] | Weight for the basis of element in x-direction at area quadrature point i | managed |
| YFAC[k,i,e] | Weight for the basis of element in x-direction at area quadrature point i | managed |
| SRFAC[k,i,e] | Weight for the basis of the source term on element at area quadrature point i | managed |
| EDGEFAC[k,i,l] | Weight for the basis of the source term on element at area quadrature point i | constant |
| EDGE_ze[k,l,e] | Numerical flux contribution to the basis of at local edge of element | managed |
| EDGE_qx[k,l,e] | Numerical flux contribution to the basis of at local edge of element | managed |
| EDGE_qy[k,l,e] | Numerical flux contribution to the basis of at local edge of element | managed |
5.2 GPU kernels
This section outlines the main subroutines and how they map to GPU parallelism. For the integration kernels, we follow the one-thread-per-element and one-thread-per-edge approach as in [Fuhry (2013)].
5.2.1 Area integration
The volume integration routine (Algorithm 1) is naturally data-parallel, as computations are done locally to each element . At a given Runge-Kutta stage irk, it computes the terms
from (3) and accumulates them into the RHS vectors. Each element is assigned to a thread identified by its global ID (in the CUDA version, this is the variable
(blockidx%x-1)*blockdim%x + threadidx%x)
which then computes the RHS contributions. This is done by looping through the quadrature points; for each quadrature point we loop through the degrees of freedom for each state variable and compute the flux terms. Recall from Eq 2 that these terms (explicitly written in terms of the state variables) are
and these are represented by the procedures Continuity_Flux, X_Flux, and Y_Flux in the code. The source terms are calculated in the same way but retrieve additional variables such as current rain intensity and wind speed.
In the OpenACC version (Listing 2), the thread-level loop is enabled through the gang-vector parallelism over the number of elements (NE), and the inner loops over quadrature points and degrees of freedom are enforced to be serial. This needs to be explicitly inserted, as by default the inner loops will also be parallelized, and in this case we know the trip counts is too short to be profitable. Importantly, the async clause is added to eliminate kernel launch overhead, and the default(present) the overhead of checking if data is already on the device. This same structure is applied for the edge integration kernel. We have observed that using these clauses speeds up the program by almost 1.5.
5.2.2 Edge integration
Edge integration routines compute the line integrals
and accumulate them to the RHS. Similar to the volume integration kernel, a natural parallelism is in the number of edges. There are multiple ways to add contributions from inter-element fluxes. One is to let each edge compute the numerical flux between the two elements, and then atomically add back the contributions to each. Atomics are necessary since two edges can share the same element.
Another approach is to first store the flux contributions to each element separately for each of its edges, i.e., we create an additional array elem_edge(E, i) which stores the contribution at edges of element . Subsequently, in each element we then sum the three contributions locally without needing atomics. We use this approach in our work as it leads to reproducible, deterministic results due to the fixed order of summation. The pseudocode for these two steps are shown in Algorithm 2 and Algorithm 3. Note that we have abstracted away the wetting/drying branching for brevity.
In Algorithm 2, each CUDA thread is mapped to an interior edge. In line 4-9, the edge number is used to obtain the IDs of the two neighboring elements as well as their local edge numbers (1-3), with the interior side denoted with and exterior . Line integration on the edge is done by looping through the edge quadrature points. First, at each point we compute the state variables (Line 14-22), use them to compute the numerical flux. Then, we add this flux contributions to both sides of the edge, indexed using the local edge number of each element (line 34-40). The subsequent elemental flux summation is shown in Algorithm 3, where each thread is mapped to an element and sums the contributions from its 3 edges.
5.2.3 Wind kernel loops
Wind and pressure calculations are done in a series of small loops, each operating on the mesh nodes independently. In the OpenACC version, we simply parallelize each loop with !$acc parallel loop. Similarly in the CUDA version, we can use the Kernel Loop Directive !$cuf kernel do <<< *, * >>> to parallelize each loop. The reason we did not also use this in the integration kernels is that it is much more limited. For example (to our knowledge), it does not allow privatizing automatic arrays as in the OpenACC case. Using the full CUDA API also allows for future optimizations, for example shared memory manipulation, in a straightforward way.
In the CUDA version, we also declare as constant several fixed-sized arrays used in wind interpolation, which results in significant speedup.
5.2.4 Data transfers for I/O
At every fixed number of time steps specified by the user, the program writes the state variables to ascii files. Thus data will have to be copied to the CPU. While it is possible to rely on managed or unified memory system to initiate the transfers, they can incur large overheads. In the CUDA version, as we traverse through the arrays to write the values, data is fetched page-by-page when page faults are triggered. In the OpenACC code, using unified memory results in the same problem, as there are enough access counts to trigger page faults. To avoid this, in both versions we use cudaMemPrefetchAsync to transfer data at once, as shown in Listing 3.
5.2.5 Memory usage of automatic arrays
In the OpenACC code, we sometimes observe runtime memory error when parallelizing loop with many automatic (private) arrays per thread. From our observation, this seems to be due to the allocation on these arrays on the GPU stack, a different behavior from using the CUDA API. We can avoid this by limiting the number of gangs to a fixed number, e.g.
!$acc parallel loop gang vector num_gangs(1024) private(<long list of arrays>).
as opposed to allowing an automatic scheduling of the gangs.
5.3 Multiple GPUs
For the OpenACC version, we also add support for a multi-GPU setup using CUDA-aware MPI which can be built on top of existing MPI routines. Namely, we modify the packing (gather) and unpacking (scatter) of ghost data to be computed the GPU. This removes data movement between host and device, and takes advantage of the higher memory bandwidth on the GPU.
6 Experimental Results
6.1 Test cases
To observe performance improvements, we run the solver on three realistic flooding scenarios due to Hurricane Harvey (2017): one on the Neches River and two on the whole Gulf of Mexico.
Flooding in Neches River (neches)
The first scenario simulates flooding in the Lower Neches River watershed during Hurricane Harvey. This setup has been used in past compound flooding studies [Wichitrnithed et al. (2024), Loveland et al. (2021)]. The mesh is comprise of 122,839 triangular elements and 62,075 nodes and is plotted in Figure 4. The upstream boundary is implemented using time-varying normal flux, and the downstream boundary a time-varying water elevation. Together they model the impact of river discharge in a compound flooding scenario.
Hurricane Harvey on a small mesh (harvey1)
This larger mesh is one of the cases in the ADCIRC testsuite and contains 575,512 elements and around 300,000 nodes. We impose a tidal (periodic) elevation boundary on the right side and use Harvey Best Track data from the National Hurricane Center HURDAT2 database.
Hurricane Harvey on a large mesh (harvey2)
This large setup uses the same data for Harvey as in harvey1, but the mesh is significantly more resolved (Figure 5). The mesh has been validated in past hindcasting studies and contains 6,675,517 elements and 3,352,598 nodes with a maximum resolution of 20 meters.
6.2 Results
We measure the wall clock time of each test case on each platform. In each case, we compare single-GPU performance to the baseline CPU MPI implementation using 144 cores, the maximum number for a single CPU node on GH200. To normalize the measurements across the different meshes and timestep size, we convert wall clock time to million elements per second (MEPS) as the metric for performance, calculated as
These values along with the speedups are listed in Table 5 and plotted in Figure 6. Both CUDA and OpenACC versions show similar performance and speedup over CPU code. We observe less speedup in the neches case (2.3-2.5) compared to the two Harvey cases (6.2-7.3) due to the absence of wind forcing (discussed below). Both Harvey cases show similar performances for all versions including CPU, demonstrating a good scalability on the mesh size. Minor differences can also arise from the mesh structure as well as runtime physics parameters.
| Case | No. of elements | 144 CPUs | CUDA | OpenACC |
| Neches river | 122,839 | 91.6 | 232 (2.5) | 210 (2.3) |
| Harvey 1 | 575,512 | 42 | 276 (6.6) | 261 (6.2) |
| Harvey 2 | 6,675,517 | 45.9 | 337 (7.3) | 320 (6.9) |
Next, we break down the runtime contributions from different subroutines, shown in Figure 7 for each dataset. Note that we only include the most expensive subroutines. The meaning of the various subroutine names are described in Table 6. Again, we do not observe significant differences between CUDA and OpenACC, and the most expensive subroutine is the interior edge integration, followed by wind forcing (in the harvey cases), and volume integration.
However, wind forcing takes the most time in the CPU code (60% and 56% for harvey1 and harvey2, respectively) which explains why we see greater speedups in the harvey cases over neches. Slope limiting is particularly expensive in the CPU version largely due to the communication ghost nodal values between neighboring elements across MPI domains.
| Name | Subroutine |
internal_edge |
Interior flux integration + flux gathering |
rhs_dg |
Volume integration |
rk_update |
Runge-kutta update loop for the next stage |
slopelim |
Slope limiter |
wetdry |
Wetting and drying scheme |
met_forcing |
Computation of wind and pressure |
Finally, strong scaling is evaluated for harvey2 (Figure 8) for 1-8 GPUs.
7 Discussion
This section discusses the performance results reported and how well GPU resources are being utilized. We focus on the most time-consuming kernels and discuss the harvey1 case in place of harvey2, as they are similar. Since there are no significant performance differences between CUDA and OpenACC, we only show the analysis for the CUDA version.
An important performance metric we report is the arithmetic intensity , defined as the ratio of the number of floating point operations (FLOPs) to the number of bytes transferred to/from the level in the memory hierarchy:
where can refer to the High Bandwidth (global) Memory (HBM), the L2 cache, and the L1 cache. This number, along with the performance in FLOPs per second, are obtained empirically from the NVIDIA Nsight Compute profiler. It is used in the Hierarchical Roofline Model [Yang et al. (2020)] to obtain performance bounds based on the relation:
which is based on some assumptions, e.g., latencies are ignored and computation perfectly overlaps with data transfer [Hager and Wellein (2010)]. In subroutines with low (like most in our code), performance is primarily bounded by data transfer time and will not reach the peak FLOP/s of the GPU. To plot the rooflines, we use the procedure described in [Yang (2020)].
7.1 Wind forcing kernels
As seen in the previous section, a significant performance gain lies in the wind processing kernels inside the met_forcing routine. While they consist of short loops, their performances do vary as seen in Figure 9. Wind kernels 3 and 4 reach reasonable fractions of effective performance bounds, while kernels 1 and 2 exhibit a much lower performance, when only looking at the DRAM / HBM roofline. This disparity can be attributed to the different types of calculation between these two groups of kernels. Kernels 3-5 only do simple arithmetic over each array element, while kernels 1-2 contain nested calls to functions that mostly perform repeated data accesses. This is indicated by a large separation between the HBM and L1/L2 intensities, showing good locality in the higher-level caches. In this case, performance is bound by the L2 bandwidth. As for wind kernel 5, performance is likely limited by the low reported occupancy of only 28% due to the use of many registers.
7.2 Integration kernels
Similar to the wind kernels, integration kernels are bandwidth-bound (Figure 10). However, we do observe good locality in the caches as indicated by the separations. In particular, the volume integration kernel reaches quite close to the L1 bandwidth ceiling, and similarly the L2 ceiling for the edge integration kernel. The main performance inhibitor, however, is likely the low occupancy (only 14% and 20% for edge and volume, respectively) resulting from the high register usage from each kernel (168 and 128 per thread). This is a hard constraint on our code, as we have a large number of variables (which are also likely to grow in the future) required by the various equations in the model. In fact, we had to use the launch_bounds kernel attribute to limit the amount of registers used (255) and increase occupancy, which was originally below 10% for both kernels.
Other memory-related performance factors include non-coalesced (strided) accesses to the state arrays global access patterns. As seen in Table 4, many arrays are arranged with the element id as the second index, which leads to accesses with a stride of 3 (the degrees of freedom) when each element is assigned to a thread. Global access pattern is also an issue for the edge integration kernel, as each edge must access neighboring elements where are not arranged contiguously in memory, a problem inherent in the use of unstructured meshes [Hadade et al. (2020)]. Finally, we also note that using linear () elements is typically insufficient to reach close to peak performance of the GPU [Fuhry (2013), Klöckner et al. (2009), Abdi et al. (2019)].
8 Conclusion and future work
We have created a GPU implementation of DG-SWEM using two different approaches: CUDA Fortran and OpenACC. In both cases, the porting is highly straightforward through the use of modern API on the NVIDIA Fortran compiler. We exploit loop-level parallelism from independent element/edge computations inherent in the DG method, and tested both implementations on the new Grace-Hopper system using realistic flood scenarios. We observe significant speedups over the original CPU-MPI implementation on a single node (144 cores) of 2.3-2.5 for a riverine flooding case, and over 6 for the storm surge cases. We also observe that the OpenACC version, when combined with the CUDA Unified Memory System (using -unified) on Grace-Hopper, requires very few code modifications and can be maintained in the same code base; it also offers a performance close to the current CUDA Fortran version. However, some care is needed in adding specific clauses to remove overhead and prevent a too fine-grained parallelization.
9 Data Availability
Some or all data, models, or code that support the findings of this study are available from the corresponding author upon reasonable request.
Flood datasets used
DG-SWEM code - https://github.com/foci/dgswem
10 Acknowledgements
This work has been supported by the United States National Science Foundation NSF PREEVENTS Track 2 Program, under NSF Grant Numbers 1855047 This material is also based on work supported by the US Department of Homeland Security under Grant No.2015-ST-061-ND0001-01. The views and conclusions contained in this document are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the US Department of Homeland Security.
The authors also would like to gratefully acknowledge the use of the ”ADCIRC” allocation on the Vista supercomputer, and “ADCIRC”, “DMS23001”, and “DMS21031” allocations on the Frontera supercomputer at the Texas Advanced Computing Center at the University of Texas at Austin.

### v2

@@
-Wichitrnithed, August 28, 2025 GPU-acceleration of the Discontinuous Galerkin Shallow Water Equations Solver (DG-SWEM) using CUDA and OpenACC
+Wichitrnithed, September 21, 2025 GPU-acceleration of the Discontinuous Galerkin Shallow Water Equations Model (DG-SWEM) with OpenACC
 Abstract
-This paper presents a porting of DG-SWEM, a discontinuous Galerkin solver for coastal ocean circulation, and in particular storm surge, to GPU using two separate approaches: CUDA Fortran and OpenACC.
-Time-explicit discontinuous Galerkin methods have been shown to exhibit a large amount of data parallelism due to the loose coupling between elements, and thus are naturally mapped to the GPU architecture.
-For each porting approach, we discuss the code design, ease of programming, and performance when running on realistic use cases.
-Specifically for the OpenACC version, we also aim to preserve maintainability within the same codebase through using Unified Memory.
-We test the codes on NVIDIA’s Grace Hopper chip and compare the GPU performance on each node to the MPI version on a single CPU node (144 cores).
+This paper presents a porting of DG-SWEM, a discontinuous Galerkin solver for hurricane storm surge, to NVIDIA GPUs.
+Time-explicit discontinuous Galerkin methods contain a large number of degrees of freedom but have been shown to exhibit a large amount of data parallelism due to the loose coupling between elements, and thus are naturally mapped to the GPU architecture.
+A previous framework in porting DG-SWEM to GPUs required converting subroutines from Fortran to C++ to be used with CUDA C++.
+By using OpenACC and Unified Memory, we simplify the porting process and maintain a single codebase for both CPU and GPU versions.
+We test the code using a large Hurricane Harvey scenario on NVIDIA’s Grace Hopper chip, and compare the GPU code’s performance on multiple H200 nodes to the CPU version on the same amount of Grace CPU nodes.
 1 Introduction
@@
 (2022)] which use OpenACC.
-This paper presents and compares the porting of DG-SWEM to GPU using two separate approaches: CUDA and OpenACC.
-The aim of the latter approach is to explore a directive-based programming model which requires less integration and maintenance efforts, as well as to take advantage of the new NVIDIA Grace-Hopper architecture [Singer (2024)] which provides a more efficient CPU-GPU data transfer.
 Lastly, we mention the recent open-source FEniCSx SWE solverby Dawson et al.
@@
 (2024)] and GPU accelerated by Pachev in [Pachev (2024)], where implicit schemes are considered.
-In the following, we introduce, discuss, and validate the proposed GPU acceleration.
 In Section 2, we introduce the governing model, i.e., the SWE and our DG discretization.
-Next, in Sections 3 and 4, we discuss the acceleration approach and the hardware setup we utilize.
-In 6, we present numerical studies of our methodology including benchmark tests and large scale hurricanes that include compound flood effects.
-Finally, in Sections 7 and 8, we discuss our results, draw conclusions, and discuss potential future research directions.
+Next, in Sections 3 and 4, we discuss the related work and the hardware setup we utilize.
+Section 5 details our implementation.
+In Section 6, we discuss multi-node performance results when running the code on a large Hurricane dataset.
+Finally, in Section 7 we draw conclusions and discuss potential future research directions.
 2 Model
@@
 3 GPU Programming approach
-As an alternative to CPU parallelism, graphics processing units (GPUs) can be viewed as a multithreaded vector processor [Hennessy and Patterson (2018)], where each unit executes the same instructions on multiple data (SIMD), but individual lanes can take different paths.
+As an alternative to purely distributed parallelism, graphics processing units (GPUs) can be viewed as a multithreaded vector processor [Hennessy and Patterson (2018)], where each unit executes the same instructions on multiple data (SIMD), but individual lanes can take different paths.
 To take advantage of a large number of such processors, the algorithm must exhibit a high degree of data parallelism, where multiple elements can be processed independently at the same time.
@@
 While the low-level nature of CUDA allows more opportunities for control and optimization by the programmer, it typically results in having to maintain a separate code base from the CPU code.
-Another approach to GPU programming is to use OpenACC [OpenACC-Standard.org (2021)] and OpenMP [Openmp.org (2021)]
+In the past, a framework has been created to generate GPU code from DG-SWEM through CUDA C++ [DuChene et al.
+(2011)].
+The workflow consisted of scripts that transform Fortran subroutines to C++ and handle variable declaration and allocation.
+The work noted that other alternatives at the time, such as using the PGI accelerator or CUDA-Fortran, were lacking in features or accessibility.
+A more modern approach to GPU programming is to use OpenACC [OpenACC-Standard.org (2021)] or OpenMP [Openmp.org (2021)]
 to write directives instead of manually managing individual GPU threads.
@@
 Of course, the programmer is still required to sufficiently expose the parallelism in the code, which often results in modifications of the CPU code as well.
-In this work, we port DG-SWEM using each approach and compare their implementations and performance.
-For the directive-based approach, we choose to use OpenACC as it currently has better support on the NVHPC compilers.
+We choose to use OpenACC in this work as it currently has better support on the NVHPC compilers.
+In this approach, we aim to mimimize integration and maintenance efforts, as well as to take advantage of the new NVIDIA Grace-Hopper architecture [Singer (2024)] which provides a more efficient CPU-GPU data transfer.
 4 Experimental setup
@@
 The original CPU code was compiled using the Intel oneAPI compiler stack.
-This work uses the NVIDIA SDK compiler stack (nvfortran, nvc) for both CPU and GPU versions, with the latter also using CUDA Fortran.
+This work uses the NVIDIA SDK compiler stack 24.7 (nvfortran, nvc, nvc++) for both CPU and GPU versions.
 Compilation flags for the CPU and GPU codes are listed in Table 3 .
-On Vista, the NVHPC version is 24.7 while on the local setup it is 23.11.
 | Version | Flags |
-| CUDA | nvfortran -r4 -O3 -cuda -Mextend -Mpreprocess - |
-| OpenACC | nvfortran -r4 -O3 -cuda -acc -Mextend -Mpreprocess -gpu=unified |
+| GPU | mpif90 -r4 -O3 -cuda -acc -Mextend -Mpreprocess -gpu=unified |
 | CPU | mpif90 -r4 -O3 -Mextend -Mpreprocess |
 5 Implementation
-This section details our implementation for both CUDA Fortran and OpenACC for different aspects of the code.
-In the Kernels section, a general parallelization is outlined, then details specific to either approaches are explained.
-5.1 Data management
+5.1 Unified Memory
 A significant part of porting code to GPUs is managing data movement.
 As a CPU (host) and a GPU (device) have separate physical memories, data must be transferred back and forth to perform the respective calculations.
-Even if some arrays are only used in device calculations, they still need to be initialized or read from files through the host, and we would have to declare device copies of those arrays.
-For example, an existing array in the code
+Traditionally, even if some arrays are only used in device calculations, they still need to be initialized or read from files through the host, and we would have to declare device copies of those arrays.
+For example, in CUDA Fortran, an existing array in the code
 real, allocatable :: A(:)
@@
 A = A_d
-To avoid having to do this for each array in the code, we use CUDA’s unified and constant memory types, explained below.
-5.1.1 Unified memory
-Unified Virtual Addressing (UVA) was introduced in CUDA 6.0.
+To avoid having to do this for each array in the code, we exploit
+Unified Memory which was introduced in CUDA 6.0.
 This allows GPU kernels to access data on the host without needing to explicitly copy the data to device.
 Data migration is automatically performed once a page fault is triggered on the device.
-One way to achieve this is through the managed attribute.
-This way, we can reuse the existing arrays in the code and only modify their declarations, e.g.
-real, allocatable, managed :: A(:,:)
-! Use A on the device
-istat = cudaDeviceSynchronize()
-! Use A on the host
-For newer systems that support Heterogenous Memory Management (HMM), the declaration can be completely eliminated using the compilation flag -gpu=unified.
-This significantly simplifies programming efforts and enhances maintainability.
-In both approaches, a synchronization is needed for correctness before the CPU can access data, since CUDA subroutines immediately return control to the host.
-In this work, the CUDA Fortran approach uses the managed declarations to make explicit to the programmer which arrays will be used on the GPU, while the OpenACC code uses -gpu=unified to avoid any data declarations.
-5.1.2 Constant memory
-The standard way of passing scalar variables to CUDA kernels is by value, i.e.
-through the Fortran value attribute in the dummy argument declaration.
-When there are a lot of variables, such as constants, this increases the programmer’s efforts to pass them correctly.
-To avoid this, we use CUDA’s constant attribute for constant variables.
-They are declared and initialized in the host global module, and reside as read-only variables on the GPU constant memory partition which a size limit of 64 KB.
-Additionally, the constant memory has faster access time than global memory when all threads access the same element — a common pattern for many fixed-size arrays like quadrature weights.
-In Table 4, we summarize the important arrays whose attributes are modified, some of which are referred to in the kernel outline in the next section.
-| Array | Meaning | Attribute |
-| ZE[k,e,irk] | mode of elevation solution at element on RK stage irk | managed |
-| QX[k,e,irk] | mode of x-velocity solution at element on RK stage irk | managed |
-| QY[k,e,irk] | mode of y-velocity solution at element on RK stage irk | managed |
-| RHS_ZE[k,e,irk] | Contribution to the mode of at element on RK stage irk | managed |
-| RHS_QX[k,e,irk] | Contribution to the mode of at element on RK stage irk | managed |
-| RHS_QY[k,e,irk] | Contribution to the mode of at element on RK stage irk | managed |
-| PHI_AREA[k,i] | basis evaluated at area quadrature point i | constant |
-| PHI_EDGE[k,i] | basis evaluated at edge quadrature point i | constant |
-| XFAC[k,i,e] | Weight for the basis of element in x-direction at area quadrature point i | managed |
-| YFAC[k,i,e] | Weight for the basis of element in x-direction at area quadrature point i | managed |
-| SRFAC[k,i,e] | Weight for the basis of the source term on element at area quadrature point i | managed |
-| EDGEFAC[k,i,l] | Weight for the basis of the source term on element at area quadrature point i | constant |
-| EDGE_ze[k,l,e] | Numerical flux contribution to the basis of at local edge of element | managed |
-| EDGE_qx[k,l,e] | Numerical flux contribution to the basis of at local edge of element | managed |
-| EDGE_qy[k,l,e] | Numerical flux contribution to the basis of at local edge of element | managed |
+In the past, this requires a special allocation, e.g.
+cudaMallocManaged in CUDA C.
+For newer systems that support Heterogenous Memory Management (HMM), all device data declarations can be completely eliminated using the compilation flag -gpu=unified.
+This significantly simplifies programming effort and enhances maintainability.
+In both approaches, a synchronization statement like !$acc wait is needed for correctness before the CPU can access the computed data, since GPU kernels immediately return control to the host.
+In Table 4 we summarize the important data arrays, some of which are referred to in the kernel descriptions in the next section.
+| Array | Meaning |
+| ZE[k,e,irk] | mode of elevation solution at element on RK stage irk |
+| QX[k,e,irk] | mode of x-velocity solution at element on RK stage irk |
+| QY[k,e,irk] | mode of y-velocity solution at element on RK stage irk |
+| RHS_ZE[k,e,irk] | Contribution to the mode of at element on RK stage irk |
+| RHS_QX[k,e,irk] | Contribution to the mode of at element on RK stage irk |
+| RHS_QY[k,e,irk] | Contribution to the mode of at element on RK stage irk |
+| PHI_AREA[k,i] | basis evaluated at area quadrature point i |
+| PHI_EDGE[k,i] | basis evaluated at edge quadrature point i |
+| XFAC[k,i,e] | Weight for the basis of element in x-direction at area quadrature point i |
+| YFAC[k,i,e] | Weight for the basis of element in x-direction at area quadrature point i |
+| SRFAC[k,i,e] | Weight for the basis of the source term on element at area quadrature point i |
+| EDGEFAC[k,i,l] | Weight for the basis of the source term on element at area quadrature point i |
+| EDGE_ze[k,l,e] | Numerical flux contribution to the basis of at local edge of element |
+| EDGE_qx[k,l,e] | Numerical flux contribution to the basis of at local edge of element |
+| EDGE_qy[k,l,e] | Numerical flux contribution to the basis of at local edge of element |
 5.2 GPU kernels
-This section outlines the main subroutines and how they map to GPU parallelism. For the integration kernels, we follow the one-thread-per-element and one-thread-per-edge approach as in [Fuhry (2013)].
+This section outlines the main subroutines and how their parallelization. For the integration kernels, we follow the one-thread-per-element and one-thread-per-edge approach as in [Fuhry (2013)].
 5.2.1 Area integration
-The volume integration routine (Algorithm 1) is naturally data-parallel, as computations are done locally to each element . At a given Runge-Kutta stage irk, it computes the terms
-from (3) and accumulates them into the RHS vectors. Each element is assigned to a thread identified by its global ID (in the CUDA version, this is the variable
-(blockidx%x-1)*blockdim%x + threadidx%x)
-which then computes the RHS contributions.
-This is done by looping through the quadrature points; for each quadrature point we loop through the degrees of freedom for each state variable and compute the flux terms.
-Recall from Eq 2 that these terms (explicitly written in terms of the state variables) are
+The area integration routine (Algorithm 1) is naturally data-parallel, as computations are done locally to each element . At a given Runge-Kutta stage irk, it computes the terms
+from (3) and accumulates them into the RHS vectors. Recall from Eq 2 that the flux terms (explicitly written in terms of the state variables) are
 and these are represented by the procedures Continuity_Flux, X_Flux, and Y_Flux in the code.
 The source terms are calculated in the same way but retrieve additional variables such as current rain intensity and wind speed.
-In the OpenACC version (Listing 2), the thread-level loop is enabled through the gang-vector parallelism over the number of elements (NE), and the inner loops over quadrature points and degrees of freedom are enforced to be serial.
-This needs to be explicitly inserted, as by default the inner loops will also be parallelized, and in this case we know the trip counts is too short to be profitable.
+There are technically three levels of loop parallelism: looping over the elements, looping over the quadrature points, and looping over the degrees of freedom.
+Here, the element-level loop is accelerated through the gang-vector parallelism over the number of elements (line 2), and the inner loops over quadrature points and degrees of freedom are enforced to be serial using !$acc seq.
+This needs to be explicitly inserted, as by default the inner loops will also be parallelized, and in this case the trip counts is too short to be profitable.
+We also add the private(...) clause to keep any local arrays private to each thread.
 Importantly, the async clause is added to eliminate kernel launch overhead, and the default(present) the overhead of checking if data is already on the device.
@@
 Note that we have abstracted away the wetting/drying branching for brevity.
-In Algorithm 2, each CUDA thread is mapped to an interior edge.
+In Algorithm 2, each GPU thread is mapped to an interior edge, again through gang-vector parallelism (line 2).
 In line 4-9, the edge number is used to obtain the IDs of the two neighboring elements as well as their local edge numbers (1-3), with the interior side denoted with and exterior .
 Line integration on the edge is done by looping through the edge quadrature points.
-First, at each point we compute the state variables (Line 14-22), use them to compute the numerical flux.
-Then, we add this flux contributions to both sides of the edge, indexed using the local edge number of each element (line 34-40).
+First, at each point we compute the state variables (Line 14-23), use them to compute the numerical flux.
+Then, we add this flux contributions to both sides of the edge, indexed using the local edge number of each element (line 35-41).
 The subsequent elemental flux summation is shown in Algorithm 3, where each thread is mapped to an element and sums the contributions from its 3 edges.
-5.2.3 Wind kernel loops
-Wind and pressure calculations are done in a series of small loops, each operating on the mesh nodes independently.
-In the OpenACC version, we simply parallelize each loop with !$acc parallel loop.
-Similarly in the CUDA version, we can use the Kernel Loop Directive !$cuf kernel do <<< *, * >>> to parallelize each loop.
-The reason we did not also use this in the integration kernels is that it is much more limited.
-For example (to our knowledge), it does not allow privatizing automatic arrays as in the OpenACC case.
-Using the full CUDA API also allows for future optimizations, for example shared memory manipulation, in a straightforward way.
-In the CUDA version, we also declare as constant several fixed-sized arrays used in wind interpolation, which results in significant speedup.
-5.2.4 Data transfers for I/O
+Note that for brevity, we have opted for the atomic approach for other less common boundary integration routines.
+5.2.3 Other kernels
+Other kernels like slope limiting (SLOPELIM) and wet/dry calculations (WETDRY) are similarly parallelized at the element-level loop, with small inner loops serialized.
+Wind forcing and pressure calculations (WIND) are done in a series of loops, each operating on the mesh nodes independently.
+We simply parallelize each loop with !$acc parallel loop.
+5.2.4 Data transfers between host and device
 At every fixed number of time steps specified by the user, the program writes the state variables to ASCII files.
 Thus data will have to be copied to the CPU.
-While it is possible to rely on managed or unified memory system to initiate the transfers, they can incur large overheads.
-In the CUDA version, as we traverse through the arrays to write the values, data is fetched page-by-page when page faults are triggered.
-In the OpenACC code, using unified memory results in the same problem, as there are enough access counts to trigger page faults.
-To avoid this, in both versions we use cudaMemPrefetchAsync to transfer data at once, as shown in Listing 3.
+While it is possible to completely rely on HMM to initiate the transfers, they can incur large overhead as there are enough access counts to trigger page migration.
+To avoid this, we use cudaMemPrefetchAsync to transfer data at once, as shown in Listing 1.
 5.2.5 Memory usage of automatic arrays
-In the OpenACC code, we sometimes observe runtime memory error when parallelizing loop with many automatic (private) arrays per thread.
+We sometimes observe runtime memory error when parallelizing loop with many automatic (private) arrays per thread.
 From our observation, this seems to be due to the allocation on these arrays on the GPU stack, a different behavior from using the CUDA API.
@@
 !$acc parallel loop gang vector num_gangs(1024) private(<long list of arrays>).
-as opposed to allowing an automatic scheduling of the gangs.
+as opposed to allowing an automatic scheduling of gangs.
 5.3 Multiple GPUs
-For the OpenACC version, we also add support for a multi-GPU setup using CUDA-aware MPI which can be built on top of existing MPI routines.
-Namely, we modify the packing (gather) and unpacking (scatter) of ghost data to be computed the GPU.
-This removes data movement between host and device, and takes advantage of the higher memory bandwidth on the GPU.
-6 Experimental Results
-6.1 Test cases
-To observe performance improvements, we run the solver on three realistic flooding scenarios due to Hurricane Harvey (2017): one on the Neches River and two on the whole Gulf of Mexico.
-Flooding in Neches River (neches)
-The first scenario simulates flooding in the Lower Neches River watershed during Hurricane Harvey.
-This setup has been used in past compound flooding studies [Wichitrnithed et al.
-(2024), Loveland et al.
-(2021)].
-The mesh is comprise of 122,839 triangular elements and 62,075 nodes and is plotted in Figure 4.
-The upstream boundary is implemented using time-varying normal flux, and the downstream boundary a time-varying water elevation.
-Together they model the impact of river discharge in a compound flooding scenario.
-Hurricane Harvey on a small mesh (harvey1)
-This larger mesh is one of the cases in the ADCIRC testsuite and contains 575,512 elements and around 300,000 nodes.
+To exchange data among different domains when using multiple processors, DG-SWEM uses MPI persistent communication.
+At the start of the simulation, calls to MPI_SEND_INIT() and MPI_RECV_INIT() are made to specify the location of data to exchange to each neighbor.
+Then, as shown in Listing 2, whenever we need to exchange data, we call MPI_STARTALL() to initiate the communication in a non-blocking manner.
+At the end, we wait until all the messages have arrived.
+The packing of the elements to send is done through the subroutine packElems(), and the unpacking similarly through unpackElems().
+The existing MPI routines need no further modification in the GPU code, as they accept unified memory addresses.
+However, we observe data being transferred to host and back to device during MPI communication.
+This appears to be a limitation of using unified memory with CUDA-aware MPI which prevents direct GPU-to-GPU data transfer [Chenbr (2025)].
+The packing and unpacking have been parallelized to take advantage of the higher memory bandwidth of the GPU, as shown in Listing 3.
+6 Results
+6.1 Test Case
+To measure performance improvements, we run the solver on a dataset for Hurricane Harvey (2017) which is representative of a typical simulation scenario.
+The mesh is shown in Figure 2 which contains 6,675,517 elements and 3,352,598 nodes with a maximum resolution of 20 meters.
 We impose a tidal (periodic) elevation boundary on the right side and use Harvey Best Track data from the National Hurricane Center HURDAT2 database.
-Hurricane Harvey on a large mesh (harvey2)
-This large setup uses the same data for Harvey as in harvey1, but the mesh is significantly more resolved (Figure 5).
-The mesh has been validated in past hindcasting studies and contains 6,675,517 elements and 3,352,598 nodes with a maximum resolution of 20 meters.
-6.2 Results
-We measure the wall clock time of each test case on each platform.
-In each case, we compare single-GPU performance to the baseline CPU MPI implementation using 144 cores, the maximum number for a single CPU node on GH200.
-To normalize the measurements across the different meshes and timestep size, we convert wall clock time to million elements per second (MEPS) as the metric for performance, calculated as
-These values along with the speedups are listed in Table 5 and plotted in Figure 6.
-Both CUDA and OpenACC versions show similar performance and speedup over CPU code.
-We observe less speedup in the neches case (2.3-2.5) compared to the two Harvey cases (6.2-7.3) due to the absence of wind forcing (discussed below).
-Both Harvey cases show similar performances for all versions including CPU, demonstrating a good scalability on the mesh size.
-Minor differences can also arise from the mesh structure as well as runtime physics parameters.
-| Case | No. of elements | 144 CPUs | CUDA | OpenACC |
-| Neches river | 122,839 | 91.6 | 232 (2.5) | 210 (2.3) |
-| Harvey 1 | 575,512 | 42 | 276 (6.6) | 261 (6.2) |
-| Harvey 2 | 6,675,517 | 45.9 | 337 (7.3) | 320 (6.9) |
-Next, we break down the runtime contributions from different subroutines, shown in Figure 7 for each dataset.
-Note that we only include the most expensive subroutines.
-The meaning of the various subroutine names are described in Table 6.
-Again, we do not observe significant differences between CUDA and OpenACC, and the most expensive subroutine is the interior edge integration, followed by wind forcing (in the harvey cases), and volume integration.
-However, wind forcing takes the most time in the CPU code (60% and 56% for harvey1 and harvey2, respectively) which explains why we see greater speedups in the harvey cases over neches.
-Slope limiting is particularly expensive in the CPU version largely due to the communication ghost nodal values between neighboring elements across MPI domains.
-| Name | Subroutine |
-internal_edge |
-Interior flux integration + flux gathering |
-rhs_dg |
-Volume integration |
-rk_update |
-Runge-kutta update loop for the next stage |
-slopelim |
-Slope limiter |
-wetdry |
-Wetting and drying scheme |
-met_forcing |
-Computation of wind and pressure |
-Finally, strong scaling is evaluated for harvey2 (Figure 8) for 1-8 GPUs.
-7 Discussion
-This section discusses the performance results reported and how well GPU resources are being utilized.
-We focus on the most time-consuming kernels and discuss the harvey1 case in place of harvey2, as they are similar.
-Since there are no significant performance differences between CUDA and OpenACC, we only show the analysis for the CUDA version.
-An important performance metric we report is the arithmetic intensity , defined as the ratio of the number of floating point operations (FLOPs) to the number of bytes transferred to/from the level in the memory hierarchy:
-where can refer to the High Bandwidth (global) Memory (HBM), the L2 cache, and the L1 cache.
-This number, along with the performance in FLOPs per second, are obtained empirically from the NVIDIA Nsight Compute profiler.
-It is used in the Hierarchical Roofline Model [Yang et al.
-(2020)] to obtain performance bounds based on the relation:
-which is based on some assumptions, e.g., latencies are ignored and computation perfectly overlaps with data transfer [Hager and Wellein (2010)].
-In subroutines with low (like most in our code), performance is primarily bounded by data transfer time and will not reach the peak FLOP/s of the GPU.
-To plot the rooflines, we use the procedure described in [Yang (2020)].
-7.1 Wind forcing kernels
-As seen in the previous section, a significant performance gain lies in the wind processing kernels inside the met_forcing routine.
-While they consist of short loops, their performances do vary as seen in Figure 9.
-Wind kernels 3 and 4 reach reasonable fractions of effective performance bounds, while kernels 1 and 2 exhibit a much lower performance, when only looking at the DRAM / HBM roofline.
-This disparity can be attributed to the different types of calculation between these two groups of kernels.
-Kernels 3-5 only do simple arithmetic over each array element, while kernels 1-2 contain nested calls to functions that mostly perform repeated data accesses.
-This is indicated by a large separation between the HBM and L1/L2 intensities, showing good locality in the higher-level caches.
-In this case, performance is bound by the L2 bandwidth.
-As for wind kernel 5, performance is likely limited by the low reported occupancy of only 28% due to the use of many registers.
-7.2 Integration kernels
-Similar to the wind kernels, integration kernels are bandwidth-bound (Figure 10).
-However, we do observe good locality in the caches as indicated by the separations.
-In particular, the volume integration kernel reaches quite close to the L1 bandwidth ceiling, and similarly the L2 ceiling for the edge integration kernel.
-The main performance inhibitor, however, is likely the low occupancy (only 14% and 20% for edge and volume, respectively) resulting from the high register usage from each kernel (168 and 128 per thread).
-This is a hard constraint on our code, as we have a large number of variables (which are also likely to grow in the future) required by the various equations in the model.
-In fact, we had to use the launch_bounds kernel attribute to limit the amount of registers used (255) and increase occupancy, which was originally below 10% for both kernels.
-Other memory-related performance factors include non-coalesced (strided) accesses to the state arrays global access patterns.
-As seen in Table 4, many arrays are arranged with the element id as the second index, which leads to accesses with a stride of 3 (the degrees of freedom) when each element is assigned to a thread.
-Global access pattern is also an issue for the edge integration kernel, as each edge must access neighboring elements where are not arranged contiguously in memory, a problem inherent in the use of unstructured meshes [Hadade et al.
-(2020)].
-Finally, we also note that using linear () elements is typically insufficient to reach close to peak performance of the GPU [Fuhry (2013), Klöckner et al.
+The timestep size is set to be s, and computation is done with single-precision.
+Figure 3 shows the performance comparison between the GPU implementation and the CPU implementation from 1 to 8 nodes, where we run the simulation for 28,800 time steps.
+As described in Section 4, each GPU node consists of a single NVIDIA H200, while each CPU node contains 144 NVIDIA Grace cores.
+Here, ”performance” is measured by the rate of work done, defined to be the number of degrees of freedom computed per second.
+In a node-by-node comparison, we observe a speedup of GPU code over CPU of around to .
+Figure 4 further shows the breakdown of time spent in each kernel.
+In this case, we only consider the total time spent in the timestepping loop (i.e.
+exclude I/O or startup time) and run the simulation for 2,880 time steps.
+Here, we see that parallel communication time in the GPU case increases with more GPUs to around 6 s at 8 nodes, resulting in a scaling plateau.
+For the same amount of nodes in the CPU case case, the communication time appears to stabilize to around 18 s.
+6.2 Performance Evaluation
+To further evaluate performance of the time-consuming kernels, we first calculate the theoretical arithmetic intensity at the DRAM level.
+This is the ratio of floating point operations (FLOPs) to total memory traffic in bytes from/to DRAM:
+We count each arithmetic operation like addition and division counts as one FLOP, including special operations like square root and power.
+In estimating from the source code (by manually counting), we also assume that everything is loaded once from memory for simplicity, which is generally not valid for large enough datasets and when accesses are not fully coalesced.
+Thus we expect the estimated values to be higher than the true values.
+The estimated arithmetic intensities for the integration kernels are shown in Table 5, along with their experimental values obtained from the NVIDIA Nsight Compute profiler.
+We focus on these two kernels as they are representative of the DG method.
+| Estimated | Experimental | |||
+| Kernel | GFLOPs | GBytes | ||
+| EDGE | 7.95 | 1.6 | 4.97 | 2.7 |
+| AREA | 4.09 | 2.2 | 1.85 | 1.81 |
+| GATHER | 0.24 | 1.92 | 0.125 | 0.09 |
+It can be seen that the arithmetic intensity of AREA is very close to the estimated value.
+This is expected, as most data are not shared between different elements, and are loaded in a (mostly) coalesced manner.
+A similar reasoning applies for GATHER.
+On the other hand, EDGE has a markedly lower intensity than estimated, a little over half.
+This can be attributed to the different memory access patterns.
+While area integration loops over elemental arrays sequentially, edge integration requires each edge to fetch data from neighboring elements which may not lie close together in memory.
+This poor spatial locality is a well-known issue in using unstructured meshes [Hadade et al.
+(2020)] and in our case results in uncoalesced accesses and therefore more data traffic.
+We further evaluate the performance using the empirical Roofline which prescribes a performance bound through the relation:
+which makes some assumptions, e.g., latencies are ignored and computation perfectly overlaps with data transfer [Hager and Wellein (2010)].
+In kernels with low (like most in our code), performance will, at best, reach only a fraction of the peak FLOP/s of the GPU.
+To plot the empirical roofline, we use the procedure described in [Yang (2020)]; this is shown in Figure 5.
+The AREA, EDGE, and GATHER kernels achieve compute throughputs of 1.9 TFLOPs/s, 1.3 TFLOPs/s, and 0.24 TFLOPs/s, respectively.
+Other relevant performance characteristics from Nsight Compute are shown in Table 6.
+A few observations can be made.
+First, the INT instruction count shows that a large portion of the computation are in fact array index calculations.
+While this is largely unavoidable, they do lower the measured FLOP rate, as we cannot assume they overlap with FP32 instructions.
+Second, the achieved occupancy for the first two kernels are quite low due to the high register usage, and this can also limit performance.
+While it is possible to tune register usage through the -maxregcount compiler flag, we found that the penalty from spilling largely offsets the speedup from higher occupancy.
+We also note that using linear () elements is typically insufficient to reach close to peak performance of the GPU [Fuhry (2013), Klöckner et al.
 (2009), Abdi et al.
 (2019)].
-8 Conclusion and future work
-We have created a GPU implementation of DG-SWEM using two different approaches: CUDA Fortran and OpenACC.
-In both cases, the porting is highly straightforward through the use of modern API on the NVIDIA Fortran compiler.
-We exploit loop-level parallelism from independent element/edge computations inherent in the DG method, and tested both implementations on the new Grace-Hopper system using realistic flood scenarios.
-We observe significant speedups over the original CPU-MPI implementation on a single node (144 cores) of 2.3-2.5 for a riverine flooding case, and over 6 for the storm surge cases.
-We also observe that the OpenACC version, when combined with the CUDA Unified Memory System (using -unified) on Grace-Hopper, requires very few code modifications and can be maintained in the same code base; it also offers a performance close to the current CUDA Fortran version.
-However, some care is needed in adding specific clauses to remove overhead and prevent a too fine-grained parallelization.
+| Kernel | Registers per thread | Occupancy (%) | FP32 Inst. | INT Instr. |
+| EDGE | 249 | 9.4 | 197 | |
+| AREA | 163 | 14.9 | ||
+| GATHER | 72 | 40.6 |
+6.3 Lines of code changed
+To get a rough estimate of the level of programming effort, we compare the total lines of code before and after incorporating OpenACC (we ignore blank lines).
+Lines of code increase by 147, from 48,674 to 48,821; this change includes the addition OpenACC directives and other modifications such as the flux-gathering routine and MPI packing/unpacking.
+Table 7 shows the breakdown of the specific OpenACC directives used.
+| Directive | Description | # of lines |
+!$acc parallel |
+Loop parallelism | 29 |
+!$acc routine |
+Convert of CPU to GPU routines | 15 |
+!$acc loop seq |
+Serialize loops | 27 |
+!$acc atomic |
+Atomic updates | 15 |
+!$acc wait |
+Barrier | 8 |
+| Continuation lines | 4 | |
+| Total | 98 |
+7 Conclusion
+We have ported DG-SWEM to NVIDIA GPUs using OpenACC Fortran in a basic manner by exploiting the data parallelism inherent in the explicit-time DG method through element and edge computations .
+We tested the code on multiple nodes on the NVIDIA Grace Hopper system using a large flooding test case and observed significant speedups over the original CPU implementation.
+When combined with the CUDA Unified Memory System, the porting requires very few code modifications and can be maintained in the same code base; this is a marked improvement from our previous porting framework [DuChene et al.
+(2011)].
+Finally, we note that there are potential performance issues that can be further investigated on top of this basic parallelization.
+First, a more detailed benchmark on multi-node scaling can be done to observe the code’s relative performance, and whether further optimizations like overlapping computation and communication will be beneficial.
+Second, we can explore alternative ways to structure edge integration to exploit GPU shared memory and improve data locality among neighboring edges.
 8 Data Availability
@@
 9 Acknowledgements
-This work has been supported by the United States National Science Foundation NSF PREEVENTS Track 2 Program, under NSF Grant Numbers 1855047 This material is also based on work supported by the US Department of Homeland Security under Grant No.2015-ST-061-ND0001-01.
+This work has been supported by the United States National Science Foundation NSF PREEVENTS Track 2 Program, under NSF Grant Numbers 1855047.
+This material is also based on work supported by the US Department of Homeland Security under Grant No.2015-ST-061-ND0001-01.
 The views and conclusions contained in this document are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the US Department of Homeland Security.